OpenStudy (anonymous):

Given point (-3,-5) and a slope of 2, write the equation in slope-intercept form.

OpenStudy (akashdeepdeb):

This is the point-slop formula of a line: y-y1 = m(x-x1) Put (x1,y1) = (-3,-5) and slope m=2 In the formula and write it in like this form The slope0intercept form = y=mx+b [b = constant/y-intercept] :)